# Memory Assisted LLM for Personalized Recommendation System

### v1

@@ Abstract.
Large language models (LLMs) have demonstrated significant potential in solving recommendation tasks. With proven capabilities in understanding user preferences, LLM personalization has emerged as a critical area for providing tailored responses to individuals. Current studies explore personalization through prompt design and fine-tuning, paving the way for further research in personalized LLMs. However, existing approaches are either costly and inefficient in capturing diverse user preferences or fail to account for timely updates to user history. To address these gaps, we propose the Memory-Assisted Personalized LLM (MAP). Through user interactions, we first create a history profile for each user, capturing their preferences, such as ratings for historical items. During recommendation, we extract relevant memory based on similarity, which is then incorporated into the prompts to enhance personalized recommendations. In our experiments, we evaluate MAP using a sequential rating prediction task under two scenarios: single domain, where memory and tasks are from the same category (e.g., movies), and cross-domain (e.g., memory from movies and recommendation tasks in books). The results show that MAP outperforms regular LLM-based recommenders that integrate user history directly through prompt design. Moreover, as user history grows, MAP’s advantage increases in both scenarios, making it more suitable for addressing successive personalized user requests.
Personalized Large Language Model, Recommendation System
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: XXXXXXXX; XXXX,XXXX; XX,XXccs: Information systems Personalizationccs: Information systems Recommender systems
@@ 1. Introduction
The recent rapid development of large language models (LLMs) has garnered considerable attention for their wide range of applications. LLMs have demonstrated remarkable zero-shot learning capabilities across various traditional machine learning tasks, including recommendation systems (RS). In the RS domain, LLMs are tasked with providing recommendations based on user-item interaction data, where the tasks are presented as prompts (Gao et al., 2023; Di Palma, 2023). However, these ”one-size-fits-all” LLMs have proven insufficient for addressing personalized individual requests (Chen et al., 2024). As a result, the personalization of LLMs has emerged as a critical area of focus, given its essential role in meeting users’ increasingly customized needs. Personalization in LLMs refers to the generation of tailored model responses by incorporating personal data, such as an individual’s historical behaviors (Tan and Jiang, 2023; Eapen and Adhithyan, 2023). By learning user preferences from these personal data, LLMs can significantly enhance their performance across fields that involve extensive individual requests, such as recommendation tasks (Tsai et al., 2024; Kang et al., 2023).
Researchers have extensively explored the implementation of personalized large language models (LLMs), focusing on two primary approaches: fine-tuning and prompt design. On one hand, personalized LLMs with fine-tuning have demonstrated advantages in various personalized tasks by adjusting LLM parameters based on users’ behavioral histories (Tan et al., 2024). However, fine-tuning faces significant challenges, particularly regarding the storage requirements associated with millions of distinct user histories that need to be fine-tuned (Salemi et al., 2023). Furthermore, compared to zero-shot in-context learning, fine-tuning is not only more computationally expensive but also prone to issues such as catastrophic forgetting of pre-trained knowledge and difficulties in aligning the model during the personalization process (Woźniak et al., 2024). On the other hand, LLM recommenders that incorporate user history via prompt templates have proven to be efficient, especially when enhanced with additional strategies such as limiting the recommendation space, filtering user histories, and applying few-shot prompt strategies (Dai et al., 2023; Wang and Lim, 2023; Salemi et al., 2023; Zhiyuli et al., 2023). Despite their efficiency, these approaches fail to account for the dynamic nature of user history. As a result, they have not adequately evaluated LLM-based recommendation systems across different sizes of user histories and exhibit inefficiencies in learning preferences from extended user histories.
@@
For demonstration and evaluation, we plan to test MAP in both single-domain and cross-domain recommendation scenarios. In the single-domain case, we will focus on movies, while in the cross-domain setting, we will use books as the recommendation target and movies as the cross-domain history. However, since the LaMP benchmark does not account for the effect of varying user history sizes on recommendation systems, we have defined a new task based on one of the benchmark’s existing tasks, LaMP-3 Personalized Product Rating (Salemi et al., 2023), to assess MAP with an increasing user history. Following prior research (Gao et al., 2023; Kang et al., 2023), we treat item rating prediction as a multi-class classification task. In this new task, we conduct rating predictions iteratively, varying only two aspects: the size of the user history and the target of the prediction task. This approach allows us to investigate whether increasing user history enhances the system’s prediction accuracy. For cross-domain item rating prediction, we follow a similar iterative process, with the prediction target remaining the same for each user in every round. The objective is to determine if MAP’s prediction accuracy for book ratings improves as more movie rating history is incorporated. Experimental results show that the MAP framework enhances recommendation system performance, with improvements becoming more significant over time. This trend holds not only in single-domain item recommendations but also in cross-domain scenarios, further demonstrating MAP’s effectiveness in long-term personalized recommendation tasks.
@@ 2. Related work
2.1. Recommendation System
Traditional RS
Traditional recommendation systems are primarily categorized into three core approaches: Collaborative Filtering (CF), which predicts item ratings based on the preferences of similar users; content-based recommendation, which compares item features to users’ preferences to suggest relevant items; and hybrid methods that combine these two techniques (Fu and Ma, 2021; Pan et al., 2008; Schafer et al., 2007). However, interest in these traditional approaches has declined, as their limitations have become more evident with the emergence of large language models (LLMs) in the recommendation field. These limitations include difficulties in adapting to new users, poor performance when handling large user history datasets, and the laborious task of obtaining item descriptions through manual web scraping (Acharya et al., 2023; Kumar and Sharma, 2016). A key reason for the shift in researchers’ focus is the demonstrated superiority of LLM-based recommendation systems. These systems have shown better performance than traditional RSs, often requiring only minimal training data to achieve competitive results (Dai et al., 2023). As a result, traditional RSs are increasingly viewed as less effective in solving modern recommendation tasks.
LLM-based RS
Recently, the application of LLM in the recommendation field has gained unprecedented attention, and these approaches can be broadly categorized into two major classes: fine-tuning LLMs for downstream recommendation tasks and direct prompting of LLMs. Fine-tuning approaches typically adjust model parameters or instructions using user history pairs, while direct prompting, such as zero-shot recommendations, generates outputs solely based on prompt inputs without altering the model (Zhang et al., 2023; Liu et al., 2023). In the fine-tuning paradigm, LLMs are able to acquire more domain-specific knowledge by optimizing their parameters and weights through task-specific datasets. This allows the model to become more specialized for particular recommendation tasks. However, fine-tuning is also widely recognized as computationally expensive and time-consuming, even when employing more efficient methods like Parameter-Efficient Fine-Tuning (PEFT), which reduces the resource requirements compared to full-model fine-tuning (Woźniak et al., 2024). On the other hand, the direct prompting paradigm for recommendation systems is more flexible and lightweight, as it bypasses the need for extensive model updates, offering a more efficient alternative for generating recommendations (Zhao et al., 2024).
@@ 2.2. Personalization in Large language Model
Personalizing LLMs remains an important yet underexplored research challenge. Early studies have demonstrated that LLMs can outperform popular recommendation algorithms when integrated with recommendation systems (Gao et al., 2023; Kim et al., 2024). Subsequent research efforts in personalizing LLMs have largely focused on designing prompts that incorporate user-item interaction histories. For example, Lyu et al. (Lyu et al., 2023) explored four prompting strategies for RS testing, showing that strategically guiding the LLM toward specific queries can enhance recommendation accuracy. Similarly, Dai et al. (Dai et al., 2023) used ChatGPT to explore its capabilities across three types of ranking tasks: point-wise ranking (predicting product ratings), pair-wise ranking (choosing between two product pairs), and list-wise ranking (ordering a list of products based on user preferences). Salemi et al. (Salemi et al., 2023) proposed the LaMP benchmark as an evaluation framework for LLM-based recommenders, building on existing natural language processing (NLP) benchmarks, which do not allow for model development to adapt to specific user needs. The LaMP benchmark includes seven personalized tasks, with three involving text classification and four involving text generation, covering tasks like tagging, rating, and headline generation. To address the limitations of most LLMs’ context lengths, the benchmark introduces Retrieval Augmentation, a method for retrieving the most relevant user behavior before prompting the LLM. This approach increases the likelihood that the LLM will accurately learn user preferences for a given item. However, as mentioned earlier, the benchmark does not allow for evaluations based on varying sizes of user histories, which limits its ability to assess the performance of personalized LLMs in different historical contexts.
@@ 4. Method: Memory Assisted LLM-based Recommendation System
In this section, we present the pipeline of the proposed Memory-Assisted LLM for Personalized Recommendation System (MAP). The core idea of memory assistance is to maintain a retrievable user profile, stored separately from the language model, functioning as a dynamic memory. This memory is continuously updated based on user interactions and retrieved to refine recommendation outputs by incorporating relevant user preferences. The primary objective is to enable recommendation systems (RSs) to generate more personalized responses tailored to each individual user. The MAP system consists of five key components: the user profile, detection module, retrieval module, update module, and a basic large language model. The user profile stores information on the user’s historical interactions and preferences. The detection module monitors user interactions to identify changes or new inputs. The retrieval module accesses relevant portions of the user profile to be used for making predictions. The update module ensures the memory is kept current by integrating new user data after each interaction. Finally, the base large language model processes the inputs and generates personalized recommendations, enhanced by the retrieved user history. The flow of this system is illustrated in Fig. 1.
@@ 4.3. Memory retrieval and Memory Assisted Recommendation
When a query involves generating a recommendation, the retrieval module first loads the user’s stored history from memory, which typically consists of a list of items, including details such as the name, description, and historical rating. However, loading the entire user history every time a recommendation is requested can be inefficient, as many entries may be irrelevant to the current query. Moreover, large language models (LLMs) have inherent limitations, including high computational costs for processing long inputs and constraints on maximum sequence length.
To address these challenges, we draw inspiration from previous studies that have shown the effectiveness of in-prompt retrieval augmentation for personalizing LLMs (Salemi et al., 2023). In our memory-assisted approach, rather than loading all user data, the retrieval module selectively extracts the most relevant data points from the user profile based on the current query. This strategy enables the system to focus on the most pertinent subset of data, improving both the efficiency of processing and the quality of the generated recommendations. By limiting the input to only the most relevant user preferences, the system enhances its ability to produce accurate, personalized recommendations without overwhelming the LLM with unnecessary data.
@@ 5.2. Baseline Model
Based on the approach in (Dai et al., 2023), we developed a baseline model that directly prompts the LLM with the unprocessed rating history, without using extracted memory. This baseline model serves as a comparison for evaluating the effectiveness of our memory-assisted approach using the GPT-3.5-turbo model.
@@ 5.3. Single-domain Recommendations
We first evaluate the single-domain recommendation task using movie rating prediction. In this scenario, each user is provided with a list of historical movie ratings reflecting their preferences, and the goal is to predict the next movie rating. For each user, the historical data starts with 1 entries, and with each iteration, additional rating history is incrementally added, growing from 1 to 19 entries.
@@
The memory-assisted model is evaluated using the widely adopted MovieLens 100k dataset (Harper and Konstan, 2015), which contains 944 users and 1,683 movies. To ensure sufficient data per user, we exclude users with fewer than 20 movie ratings and limit each user to exactly 20 ratings. After preprocessing, the dataset consists of 834 users and a total of 16,680 movie ratings. For evaluation, we compare the predicted rating for each user’s next movie with the corresponding ground truth rating.
@@ Results
For each user, the model generates 19 predictions for 19 different movies. We compute the absolute difference between the predictions and the users’ actual ratings and calculate the Mean Absolute Error (MAE) across the 834 users. From the visualization in Fig. 6, the memory-assisted model consistently outperforms the baseline starting from the second iteration, with similar performance to the baseline during the first iteration.
@@
Table 1. The comparison of Mean Absolute Error (MAE) of single-domain recommendation under different user history sizes
@@
Beyond performance, we also compare the cost of running the recommendation models using the same API of the large language model (see Table 2). For both our model and the baseline model, we use the pricing of GPT-3.5-turbo (OpenAI Pricing, 2024). For the MovieLens dataset, we calculate the average number of tokens processed per user history and multiply this by the API’s cost per token. The baseline model (Vanilla GPT) requires more tokens to generate recommendations due to its less efficient handling of user history, resulting in a higher operational cost of \$$\$0.00325$ per 10 histories per user. In contrast, our proposed MAP model optimizes token usage by summarizing and retrieving relevant user history, thereby reducing the cost to \$$\$0.00086$ per 10 histories per user. This significant cost reduction demonstrates the economic advantage of our approach, making it more scalable and sustainable for large-scale deployments.
@@ 5.4. Cross-domain Recommendation
To evaluate the model’s ability to perform cross-domain recommendations, we designed an experiment where the task is to predict book ratings based on a user’s movie rating history. This experiment tests the model’s ability to transfer knowledge from one domain (movies) to another (books) and still provide relevant recommendations. For each user, the experiment is conducted over 19 iterations, with the number of movie ratings used in each iteration incrementally increasing from 1 to 19. The memory-assisted model retrieves relevant movie ratings from the user’s profile and calculates the cosine similarity between movie genres and book genres. The most similar movie ratings are used to assist in predicting the user’s book rating.
@@
For this cross-domain experiment, we use movie and book ratings from the Amazon Review Data (Ni et al., 2019). We employ the ”ratings only” version of the dataset and merge it with item metadata to ensure both movie and book ratings are aligned with their respective genres. After filtering out users with fewer than 20 movie interactions and capping the number of movie ratings at 19 per user, the refined dataset consists of 418 users, 418 book ratings, and 7,942 movie ratings.
@@ Results
Table 3. The comparison of Mean Absolute Error (MAE) of cross-domain recommendation under different user history sizes.
@@
For each user, the model generates 19 predictions for a single book, each based on an incrementally growing history of movie ratings ranging from 1 to 19 entries. The absolute difference between each predicted book rating and the actual rating is calculated, and the Mean Absolute Error (MAE) is computed across all 418 users. To ensure robustness, we shuffle the order of the movie history for each user and repeat the experiment for both the memory-assisted model and the baseline model. The final MAE is averaged across the shuffled and unshuffled results to eliminate potential biases from the order of history inputs.

### v2

@@ Abstract.
Large language models (LLMs) have demonstrated significant potential in solving recommendation tasks. With proven capabilities in understanding user preferences, LLM personalization has emerged as a critical area for providing tailored responses to individuals. Current studies explore personalization through prompt design and fine-tuning, paving the way for further research in personalized LLMs. However, existing approaches are either costly and inefficient in capturing diverse user preferences or fail to account for timely updates to user history. To address these gaps, we propose the Memory-Assisted Personalized LLM (MAP). Through user interactions, we first create a history profile for each user, capturing their preferences, such as ratings for historical items. During recommendation, we extract relevant memory based on similarity, which is then incorporated into the prompts to enhance personalized recommendations. In our experiments, we define a new task that enables testing with varying memory size under two scenarios: single domain where memory and tasks are from the same category and cross-domain (e.g. memory from movies and recommendation tasks in books). The results show that MAP outperforms regular LLM-based recommenders that integrate user history directly through prompt design. Moreover, as user history grows, MAP’s advantage increases in both scenarios, making it more suitable for addressing successive personalized user requests.
@@ 1. Introduction
LLMs exhibit strong zero-shot learning abilities in traditional machine learning tasks, including recommendation systems (RS). In RS, LLMs generate recommendations from user-item interaction data framed as prompts (Gao et al., 2023; Di Palma, 2023). However, these ”one-size-fits-all” LLMs have proven insufficient for addressing personalized individual requests (Chen et al., 2024). As a result, there is growing interest in personalizing LLMs using personal data like behavioral histories to generate tailored responses (Tan and Jiang, 2023; Eapen and Adhithyan, 2023). By learning user preferences, LLMs can enhance performance in RS tasks to meet the increasingly customized needs (Tsai et al., 2024; Kang et al., 2023).
Researchers have explored personalized LLMs through two main approaches: fine-tuning and prompt design. Fine-tuning methods adjust LLM parameters based on user histories (Tan et al., 2024), but face significant challenges including massive storage requirements for millions of user histories as training data and computational costs during fine-tuning (Salemi et al., 2023), along with issues like catastrophic forgetting (Woźniak et al., 2024). Prompt-based approaches incorporating user history have proven more efficient, particularly when enhanced with strategies like history filtering and few-shot prompting (Dai et al., 2023; Wang and Lim, 2023; Salemi et al., 2023; Zhiyuli et al., 2023). However, these methods fail to account for dynamic user history sizes and exhibit inefficiencies when learning from extended user histories.
@@
For evaluation, we plan to test MAP in both single-domain and cross-domain recommendation scenarios. In the single-domain case, we will focus on movies, while in the cross-domain setting, we will use books as the recommendation target and movies as the cross-domain history. However, since existing benchmarks (e.g., LaMP) does not account for the effect of varying user history sizes on recommendation systems, we have defined a new task based on one of LaMP benchmark’s existing tasks, LaMP-3 Personalized Product Rating (Salemi et al., 2023), to assess MAP with an increasing user history. Following prior research (Gao et al., 2023; Kang et al., 2023), we treat item rating prediction as a multi-class classification task. In this new task, we conduct rating predictions iteratively, varying only two aspects: the size of the user history and the target of the prediction task. This approach allows us to investigate whether increasing user history enhances the system’s prediction accuracy. For cross-domain item rating prediction, we follow a similar iterative process, with the prediction target remaining the same for each user in every round. The objective is to determine if MAP’s prediction accuracy for book ratings improves as more movie rating history is incorporated. Experimental results show that the MAP framework enhances recommendation system performance, with improvements becoming more significant over time. This trend holds not only in single-domain item recommendations but also in cross-domain scenarios, further demonstrating MAP’s effectiveness in long-term personalized recommendation tasks.
The main contributions of our work are as follows:
  • We propose a new method named MAP that dynamically integrates user interaction history to complete recommendation tasks with efficiency.
  • We develop a new task from a popular benchmark LaMP (Salemi et al., 2023) to enable evaluation with varying user history size and historical contexts.
  • Our experiments show that MAP enhances RS performance in both single- and cross-domain scenarios, with greater improvements under growing memory size (e.g., in single domain, with 4 user history, MAP outruns the baseline for 0.039 mean absolute error (MAE); with 16 user history, MAP outruns the baseline for 0.104 MAE).
@@ 2. Related work
2.1. Fine-tuning and Prompting
After early studies have shown that LLMs can outperform popular recommendation algorithms, current work on LLM-powered RS is mainly from fine-tuning and prompting paradigms (Gao et al., 2023; Kim et al., 2024; Zhao et al., 2024). However, fine-tuning is also widely recognized as computationally expensive and time-consuming, even when employing more efficient methods like Parameter-Efficient Fine-Tuning (PEFT) (Woźniak et al., 2024). On the other hand, the direct prompting paradigm for recommendation systems is more flexible and lightweight, as it bypasses the need for extensive model updates, offering a more efficient alternative for generating recommendations (Zhao et al., 2024).
@@ 2.2. Personalization in Large language Model
Current LLM personalization research primarily focuses on incorporating user-interaction histories through prompt design (Lyu et al., 2023; Dai et al., 2023) and domain-specific knowledge integration (Yao et al., 2023; Lian et al., 2024). To improve efficiency, studies have explored long-term memory mechanisms, such as entity-based memory banks that store user attitudes from dialogue sessions (Xi et al., 2024b). However, these approaches lack relevant retrieval of user data, leading to low efficiency with long history contexts.
Retrieval Augmentation methods address this by filtering user behavior data before prompting (Salemi et al., 2024), evaluated using the LaMP benchmark with seven personalized tasks (Salemi et al., 2023). However, LaMP lacks evaluation across varying user history sizes, limiting assessment of personalized LLMs in different contexts. Recent frameworks have introduced optimization techniques like sequential sentiment labeling, user embedding generation, and recommendation simulation (Lee and Kim, 2025; Ning et al., 2025; Xi et al., 2024a; Zhang et al., 2024), but these require costly fine-tuning compared to direct prompting approaches, use offline datasets for implementation, or result in an unscalable memory units.
In this work, the proposed MAP furthers the study of personalized LLM-powered RS in user and item indexing and data augmentation according to suggestions of previous works on future directions (Zhao et al., 2024).
@@ 4. Method: Memory Assisted LLM-based Recommendation System
In this section, we present the pipeline of the proposed Memory-Assisted LLM for Personalized Recommendation System (MAP). The flow of this system is illustrated in Fig. 1. The core idea of memory assistance is to maintain a retrievable user profile, stored separately from the language model, functioning as a dynamic memory. This memory is continuously updated based on user interactions and retrieved to refine recommendation outputs by incorporating relevant user preferences. The primary objective is to enable recommendation systems (RSs) to generate more personalized responses tailored to each individual user. The MAP system consists of five key components: the user profile, detection module, retrieval module, update module, and a basic large language model. The user profile stores information on the user’s historical interactions and preferences. The detection module monitors user interactions to identify changes or new inputs. The retrieval module accesses relevant portions of the user profile to be used for predictions. The update module ensures the memory is kept current by integrating new data after each interaction. Finally, the base large language model processes the inputs and generates personalized recommendations, enhanced by the retrieved user history.
@@ 4.3. Memory retrieval and Memory Assisted Recommendation
When a query involves generating a recommendation, the retrieve module processes the recommendation request by augmenting it with filted memory from the user profile. The work flow of the retrieve module is illustrated in Fig. 3.
Retrieve User History
The module first loads the user’s stored history from memory, which typically consists of a list of items, including details such as the name, description, and historical rating. However, loading the entire user history every time a recommendation is requested can be inefficient, as many entries may be irrelevant to the current query. Moreover, large language models (LLMs) have inherent limitations, including high computational costs for processing long inputs and constraints on the maximum sequence length. To address these challenges, we draw inspiration from previous studies that have shown the effectiveness of in-prompt retrieval augmentation for personalizing LLMs (Salemi et al., 2023). In our memory-assisted approach, rather than loading all user data, the retrieval module selectively extracts the most relevant data points from the user profile based on the current query. This strategy enables the system to focus on the most pertinent subset of data, improving both the efficiency of processing and the quality of the generated recommendations. By limiting the input to only the most relevant user preferences, the system enhances its ability to produce accurate, personalized recommendations without overwhelming the LLM with unnecessary data.
Calculate Similarity Score & Form Memory
@@ 5.2. Baseline Model
Based on previous approach (Dai et al., 2023), we developed a baseline model that directly prompts the LLM with the unprocessed rating history, without using extracted memory. This baseline model serves as a comparison for evaluating the effectiveness of our memory-assisted approach using the GPT-3.5-turbo model.
@@ 5.3. Single-domain Recommendations
We first evaluate the single-domain recommendation task using movie rating prediction. In this scenario, each user is provided with a list of historical movie ratings reflecting their preferences, and the goal is to predict the next movie rating. For each user, the historical data starts with 1 entries, and with each iteration, additional rating history is incrementally added, growing from 1 to 18 entries.
@@
The memory-assisted model is evaluated using the widely adopted MovieLens 100k dataset (Harper and Konstan, 2015), which contains 944 users and 1,683 movies. To ensure sufficient data per user, we exclude users with fewer than 19 movie ratings and limit each user to exactly 19 ratings. After preprocessing, the dataset consists of 839 users and a total of 15,941 movie ratings. For evaluation, we compare the predicted rating for each user’s next movie with the corresponding ground truth rating.
@@ Results
For each user, the model generates 18 predictions for 18 different movies. We compute the absolute difference between the predictions and the users’ actual ratings and calculate the Mean Absolute Error (MAE) across the 834 users. From the visualization in Fig. 6, the memory-assisted model consistently outperforms the baseline starting from the second iteration, with similar performance to the baseline during the first iteration.
@@
Table 1. Mean Absolute Error (MAE) of single-domain recommendation under different user history sizes
@@
Beyond performance, we also compare the cost of running the recommendation models using the same API of the large language model (see Table 2). For both our model and the baseline model, we use the pricing of GPT-3.5-turbo (OpenAI Pricing, 2024). For the MovieLens dataset, we calculate the average number of tokens processed per user history and multiply this by the API’s cost per token. The baseline model (Vanilla GPT) requires more tokens to generate recommendations due to its less efficient handling of user history, resulting in a higher operational cost of $\mathdollar 0.00325$ per 10 histories per user. In contrast, our proposed MAP model optimizes token usage by summarizing and retrieving relevant user history, thereby reducing the cost to $\mathdollar 0.00086$ per 10 histories per user. This significant cost reduction demonstrates the economic advantage of our approach, making it more scalable and sustainable for large-scale deployments.
@@ 5.4. Cross-domain Recommendation
To evaluate the model’s ability to perform cross-domain recommendations, we designed an experiment where the task is to predict book ratings based on a user’s movie rating history. This experiment tests the model’s ability to transfer knowledge from one domain (movies) to another (books) and still provide relevant recommendations. For each user, the experiment is conducted over 18 iterations, with the number of movie ratings used in each iteration incrementally increasing from 1 to 18. The memory-assisted model retrieves relevant movie ratings from the user’s profile and calculates the cosine similarity between movie genres and book genres. The most similar movie ratings are used to assist in predicting the user’s book rating.
@@
For this cross-domain experiment, we use movie and book ratings from the Amazon Review Data (Ni et al., 2019). We employ the ”ratings only” version of the dataset and merge it with item metadata to ensure both movie and book ratings are aligned with their respective genres. After filtering out users with fewer than 18 movie interactions and capping the number of movie ratings at 18 per user, the refined dataset consists of 533 users, 533 book ratings, and 9,594 movie ratings.
@@ Results
Table 3. Mean Absolute Error (MAE) of cross-domain recommendation under different user history sizes.
@@
For each user, the model generates 18 predictions for a single book, each based on an incrementally growing history of movie ratings ranging from 1 to 18 entries. The absolute difference between each predicted book rating and the actual rating is calculated, and the Mean Absolute Error (MAE) is computed across all users. To ensure robustness, we shuffle the order of the movie history for each user and repeat the experiment for both the memory-assisted model and the baseline model. The final MAE is averaged across the shuffled and unshuffled results to eliminate potential biases from the order of history inputs.